### v1

# New Borwein-type conjectures

Abstract.
Motivated by recent research of Krattenthaler and Wang, we propose five new “Borwein-type” conjectures modulo $3$ and two new “Borwein-type” conejctures modulo $5$.
Key words and phrases:
Borwein conjectures, Borwein-type conjectures, sign-pattern.
2020 Mathematics Subject Classification:
05A16, 05A17, 05A30
1. Introduction
In $1990$, Peter Borwein unearthed several mysteries involving sign patterns of coefficients of certain polynomials. He then personally communicated George Andrews who stated them as the following in his paper [1]:
Conjecture 1.1.
(First Borwein Conjecture) For the polynomials $A_{n}(q)$, $B_{n}(q)$, and $C_{n}(q)$ defined by
$\prod\limits_{j=1}^{n}(1-q^{3j-2})(1-q^{3j-1})=A_{n}(q^{3})-qB_{n}(q^{3})-q^{2}C_{n}(q^{3}),$
each has non-negative coefficients.
Conjecture 1.2.
(Second Borwein Conjecture) For the polynomials $\alpha_{n}(q)$, $\beta_{n}(q)$, and $\gamma_{n}(q)$ defined by
$\prod\limits_{j=1}^{n}(1-q^{3j-2})^{2}(1-q^{3j-1})^{2}=\alpha_{n}(q^{3})-q\beta_{n}(q^{3})-q^{2}\gamma_{n}(q^{3}),$
each has non-negative coefficients.
Conjecture 1.3.
(Third Borwein Conjecture) For the polynomials $\nu_{n}(q)$, $\phi_{n}(q)$, $\chi_{n}(q)$, $\psi_{n}(q)$, and $\omega_{n}(q)$ defined by
$\prod\limits_{j=1}^{n}(1-q^{5j-4})(1-q^{5j-3})(1-q^{5j-2})(1-q^{5j-1})$
$=\nu_{n}(q^{5})-q\phi_{n}(q^{5})-q^{2}\chi_{n}(q^{5})-q^{3}\psi(q^{5})-q^{4}\omega(q^{5}),$
each has non-negative coefficients.
Although Andrews did not prove any of the above conjectures, he stated [1, eq. $(1.2)$] a partition theoretic analogue of Conjecture 1.1 and also derived explicit sum representations of the polynomials $A_{n}(q)$, $B_{n}(q)$, and $C_{n}(q)$. It was recently proved analytically by Chen Wang using saddle point method [2]. Apart from other tools, his proof relied on a theorem of Andrews [1, Theorem $4.1$] and the following three recursive relations also given by Andrews [1, eq. $(3.1)$ $(3.3)$]:
$A_{n}(q)$ $=(1+q^{2n-1})A_{n-1}(q)+q^{n}B_{n-1}(q)+q^{n}C_{n-1}(q),$
$B_{n}(q)$ $=(1+q^{2n-1})B_{n-1}(q)+q^{n-1}A_{n-1}(q)-q^{n}C_{n-1}(q),$
$C_{n}(q)$ $=(1+q^{2n-1})C_{n-1}(q)+q^{n-1}A_{n-1}(q)-q^{n-1}B_{n-1}(q).$
Recently, Christian Krattenthaler and Chen Wang [3] proved Conjecture 1.1 and Conjecture 1.2 completely. They also gave “two thirds” proof of the following cubic version which was missed by both Borwein and Andrews and can be stated as follows:
Conjecture 1.4.
(Cubic Borwein Conjecture) For the polynomials $\kappa_{n}(q)$, $\delta_{n}(q)$, and $\theta_{n}(q)$ defined by
$\prod\limits_{j=1}^{n}(1-q^{3j-2})^{3}(1-q^{3j-1})^{3}=\kappa_{n}(q^{3})-q\delta_{n}(q^{3})-q^{2}\theta_{n}(q^{3}),$
each has non-negative coefficients.
Unlike Wang [2] who considered Andrews’ sum representations for his proof, Krattenthaler and Wang [3] applied the saddle point method directly to the “First Borwein polynomial” $\prod\limits_{j=1}^{n}(1-q^{3j-2})(1-q^{3j-1})$ and its higher powers. They also provided ideas [3, Conjecture $11.1$] for (at least) “three fifths” proof of Conjecture 1.3 when the modulus $3$ is replaced by $5$.
From here on, let us consider the standard notation for $q$-Pochammer symbols,
$(a;q)_{n}$ $=\prod\limits_{k=0}^{n-1}(1-aq^{k}),\,\,\text{for $n\geq 1$},$
$(a;q)_{0}$ $=1.$
For positive integers $k$ and $i$, we consider polynomials of the form
$P_{n,i}^{(k)}(q):=\dfrac{(q;q)_{kn}^{i}}{(q^{k};q^{k})_{n}^{i}}.$
For $k=7$ above, Krattenthaler and Wang [3, Conjecture $11.5$] stated an interesting modulus $7$ Borwein Conjecture which is as follows:
Conjecture 1.5.
(Modulus $7$ Borwein Conjecture) For positive integers $n$, consider the expansion of the polynomial
$P_{n,1}^{(7)}(q):=\dfrac{(q;q)_{7n}}{(q^{7};q^{7})_{n}}=\sum\limits_{m=0}^{21n^{2}}d_{m}(n)q^{m}.$
Then
$d_{7m}(n)\geq 0\quad\,\text{and}\quad\,d_{7m+1}(n),d_{7m+3}(n),d_{7m+4}(n),d_{7m+6}(n)\leq 0,\quad\,\text{for all $m$ and $n$},$
while
$d_{7m+5}(n)\begin{cases}\geq 0,\quad\text{for}\,\,\,m\leq 3\alpha(n)n^{2},\\ \leq 0,\quad\text{for}\,\,\,m>3\alpha(n)n^{2},\\ \end{cases}$
where $\alpha(n)$ seems to stabilise around $0.302$.
Motivated by Conjecture 1.5, in Section 2, we state new “Borwein-type” sign-pattern conjectures for the polynomial $P_{n,i}^{(k)}(q)$ where $(k,i)\in\{(3,4),(3,5),(3,6),(3,7),(3,8),(5,2),(5,3)\}$. We conclude with some possible future directions in Section 3.
2. New Conjectures
In this section, we state the new “Borwein-type” conjectures.
Conjecture 2.1.
Let $n$ be a positive integer and $i\in\{4,5,6,7,8\}$. Consider the expansion of the polynomial $P_{n,i}^{(3)}(q)$ defined by
$P_{n,i}^{(3)}(q):=\dfrac{(q;q)_{3n}^{i}}{(q^{3};q^{3})_{n}^{i}}=\sum\limits_{m=0}^{3in^{2}}c_{m}^{(i)}(n)q^{m}.$
Then
$c_{3m}^{(i)}(n)\geq 0,\quad\,\text{for all $m$ and $n$},$
while
$c_{3m+2}^{(i)}(n)\begin{cases}\geq 0,\quad\text{for}\,\,\,m\leq\alpha^{(i)}(n)n^{2},\\ \leq 0,\quad\text{for}\,\,\,m>\alpha^{(i)}(n)n^{2},\\ \end{cases}$
where $\alpha^{(i)}(n)$ seems to stabilise around according to the following table
value of $i$ limiting value of $\alpha^{(i)}(n)$
$4$ $0.751$
$5$ $1.270$
$6$ $1.778$
$7$ $2.282$
$8$ $2.784$
Remark 1.
Since the polynomial $P_{n,i}^{(3)}(q)$ is palindromic for any positive integer $i$, we can also predict the signs of the coefficients $c_{3m+1}^{(i)}(n)$ in Conjecture 2.1. This is true because of the following duality result which is easy to prove.
Theorem 2.2.
Consider the polynomial $P_{n,i}^{(3)}(q)$ as defined in Conjecture 2.1. If we write
$P_{n,i}^{(3)}(q)=X_{n,i}(q^{3})+qY_{n,i}(q^{3})+q^{2}Z_{n,i}(q^{3}),$
then the following relations hold
$X_{n,i}(q)=q^{in^{2}}X_{n,i}(1/q)$
and
$Z_{n,i}(q)=q^{in^{2}-1}Y_{n,i}(1/q).$
Conjecture 2.3.
Let $n$ be a positive integer and $i\in\{2,3\}$. Consider the expansion of the polynomial $P_{n,i}^{(5)}(q)$ defined by
$P_{n,i}^{(5)}(q):=\dfrac{(q;q)_{5n}^{i}}{(q^{5};q^{5})_{n}^{i}}=\sum\limits_{m=0}^{10in^{2}}d_{m}^{(i)}(n)q^{m}.$
Then
$d_{5m}^{(i)}(n)\geq 0,\quad\,\text{for all $m$ and $n$},$
while
$d_{5m+3}^{(i)}(n)\begin{cases}\geq 0,\quad\text{for}\,\,\,m\leq\tilde{\alpha}^{(i)}(n)n^{2},\\ \leq 0,\quad\text{for}\,\,\,m>\tilde{\alpha}^{(i)}(n)n^{2},\\ \end{cases}$
and
$d_{5m+4}^{(i)}(n)\begin{cases}\geq 0,\quad\text{for}\,\,\,m\leq\tilde{\beta}^{(i)}(n)n^{2},\\ \leq 0,\quad\text{for}\,\,\,m>\tilde{\beta}^{(i)}(n)n^{2},\\ \end{cases}$
where $\tilde{\alpha}^{(i)}(n)$ and $\tilde{\beta}^{(i)}(n)$ seem to stabilise around according to the following table
value of $i$ limiting value of $\tilde{\alpha}^{(i)}(n)$ limiting value of $\tilde{\beta}^{(i)}(n)$
$2$ $1.133$ $1.011$
$3$ $2.132$ $2.001$
Remark 2.
Since the polynomial $P_{n,i}^{(5)}(q)$ is palindromic for any positive integer $i$, we can also predict the signs of the coefficients $d_{5m+1}^{(i)}(n)$ and $d_{5m+2}^{(i)}(n)$ in Conjecture 2.3. This is true because of the following duality result which we leave as an exercise for the interested reader to prove.
Theorem 2.4.
Consider the polynomial $P_{n,i}^{(5)}(q)$ as defined in Conjecture 2.3. If we write
$P_{n,i}^{(5)}(q)=A_{n,i}(q^{5})+qB_{n,i}(q^{5})+q^{2}C_{n,i}(q^{5})+q^{3}D_{n,i}(q^{5})+q^{4}E_{n,i}(q^{5}),$
then the following relations hold
$A_{n,i}(q)=q^{2in^{2}}A_{n,i}(1/q),$
$D_{n,i}(q)=q^{2in^{2}-1}C_{n,i}(1/q),$
and
$E_{n,i}(q)=q^{2in^{2}-1}B_{n,i}(1/q).$
Remark 3.
It must be noted that Krattenthaler and Wang [3, Remark $11.6.(2)$] mention that for $\delta\in\{2,3\}$, the coefficient of $q^{5m}$ in $(q;q)_{5n}^{\delta}/(q^{5};q^{5})_{n}^{\delta}$ is non-negative for all $m$, however, they don’t find the limiting values of $\tilde{\alpha}^{(\delta)}(n)$ and $\tilde{\beta}^{(\delta)}(n)$.
3. Concluding Remarks
  1. (1)
    It would be very interesting to investigate whether the asymptotic approach of Krattenthaler and Wang in [3] applies to Conjecture 2.1 and Conjecture 2.3.
  2. (2)
    The sign-pattern of coefficients is much more complicated for $P_{n,i}^{(3)}(q)$ when $i\geq 9$ and for $P_{n,i}^{(5)}(q)$ when $i\geq 4$ but is definitely something worth exploring more.
4. Acknowledgments
[…]